Sweeten Company had no jobs in progress at the beginning of March and no beginning inventories. The company has two manufacturing departments--Molding and Fabrication. It started, completed, and sold only two jobs during March—Job P and Job Q. The following additional information is available for the company as a whole and for Jobs P and Q (all data and questions relate to the month of March):

Molding Fabrication Total
Estimated total machine-hours used 2,500 1,500 4,000
Estimated total fixed manufacturing overhead $ 12,000 $ 16,200 $ 28,200
Estimated variable manufacturing overhead per machine-hour $ 2.20 $ 3.00
Job P Job Q
Direct materials $ 21,000 $ 12,000
Direct labor cost $ 27,400 $ 10,700
Actual machine-hours used:
Molding 2,500 1,600
Fabrication 1,400 1,700
Total 3,900 3,300

Sweeten Company had no underapplied or overapplied manufacturing overhead costs during the month.assume that Sweeten Company uses a plantwide predetermined overhead rate with machine-hours as the allocation base

5. What was the total manufacturing cost assigned to Job Q? (Do not round intermediate calculations.)

-If Job Q included 30 units, what was its unit product cost? (Do not round intermediate calculations. Round your final answer to nearest whole dollar.)

-

7. Assume that Sweeten Company used cost-plus pricing (and a markup percentage of 80% of total manufacturing cost) to establish selling prices for all of its jobs. What selling price would the company have established for Jobs P and Q? What are the selling prices for both jobs when stated on a per unit basis assuming 20 units were produced for Job P and 30 units were produced for Job Q? (Do not round intermediate calculations. Round your final answers to nearest whole dollar.)

- What was Sweeten Company’s cost of goods sold for March? (Do not round intermediate calculations.)
Answer #1

Plantwide predetermine overhead rate :

Estimated manufacturing overhead = (2500*2.2+1500*3+28200) = $38200

Predetermine overhead rate = 38200/4000 = $9.55 per machine hour

Calculate manufacturing overhead applied

Job Q
Overhead applied 3300*9.55 = $31515

5) Calculate total manufacturing cost applied job Q :

Job Q
Direct material 12000
Direct labor 10700
Overhead applied 31515
Total manufacturing cost 54215

6) Unit product cost = 54215/30 = $1807 per unit

7) Calculate Selling price :

Job P Job Q
Total manufacturing cost 85645 54215
Markup @80% 68516 43372
Selling price of job 154161 97587
Selling price per unit 154161/20 = 7708 97587/30 = 3253

Cost of goods sold for march = 154161+97587 = $251748